OpenStudy (anonymous):

2x+y+z= - 2 3x-2y-2z= -3 -5x + 3y+4z=7 Solve the system algebraically and check.. someone can help me here i will give medals and im gonna be your fan

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@shinalcantara

OpenStudy (anonymous):

can anyone can help me :D

OpenStudy (anonymous):

please help me guys.. we will have an exam tommorow i need to solve this

OpenStudy (anonymous):

largely donkey work, i would get a computer to do it http://www.wolframalpha.com/input/?i=2x%2By%2Bz%3D-2+and+3x-2y-2z%3D-3++and+-5x+%2B+3y%2B4z%3D7

OpenStudy (anonymous):

but if course this will not help you on a test lets see if we can eliminate one of the variables \[3x-2y-2z= -3 \\-5x + 3y+4z=7\] multiply the first one by 2 and get \[6x-4y-4z= -6\\ -5x + 3y+4z=7\] which after adding gets \[x-y=1\]

OpenStudy (anonymous):

repeat with \[2x+y+z= - 2\\ 3x-2y-2z= -3\][ again multiply the first one by 2 and get \[4x+2y+2z= - 4\\ 3x-2y-2z= -3\] adding gets \[7x=-7\] making \(x=1\)

OpenStudy (anonymous):

ok that was wrong \[7x=-7\\ x=-1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now that you know \(x=-1\) and \(x-y=1\) that tells you \(y=-2\)

OpenStudy (anonymous):

replace \(x\)by \(-1\) and \(y\) by \(-2\) in any of the three equations to find that \(z=2\)

OpenStudy (shinalcantara):

2x + y + z= -2 ---equation 1 3x - 2y - 2z= -3 ---equation 2 -5x + 3y + 4z=7 ---equation 3 ----------------- For this system, you can do Substitution method ----------------- From equation 1 2x + y + z = -2 y = -2x - z - 2 ----equation 1' ----------------- From equation 2 3x - 2y - 2z = -3 2y = 3x - 2z + 3 y = 1/2 (3x - 2z + 3) ----equation 2' ----------------- Equate equations 1' and 2' -2x - z - 2 = 1/2 (3x - 2z + 3) Cross multiply 2 -4x - 2z - 4 = 3x - 2z + 3 -4x - 3x - 2z + 2z = 3 + 4 -7x = 7 x = -1 ------------------ From equation 3 -5x + 3y + 4z=7 3y = 5x - 4z + 7 y = 1/3[5x - 4z + 7] ----equation 3' Substitute x=-1 y = 1/3[5(-1) - 4z + 7] y = 1/3 [-5 - 4z + 7] y = 1/3 [-4z + 2] ----equation 4 ----------------- From equations 1' y = -2x - z - 2 (x=-1) y = -2(-1) - z - 2 y = 2 - z - 2 y = - z -------equation 5 Equate equation 5 to 4 -z = 1/3 [-4z + 2] -3z = -4z + 2 -3z + 4z = 2 z = 2 ------------- Substitute z=2 to equation 5 y = -z y = -2